OpenStudy (anonymous):

Evaluate 27 × a-2 for a = 3

OpenStudy (anonymous):

is that negative 2 or minus 2

OpenStudy (lgbasallote):

\[\huge 27 \times a - 2\] when a = 3 \[\huge 27 \times 3 - 2\] where are you getting confused?

OpenStudy (anonymous):

-162

OpenStudy (anonymous):

if thats a negative 2 but if its minus 2 then its 79

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

can i just ask where -162 came from? 27 x 3 x -2?

OpenStudy (anonymous):

27*3(-2)= -162 27*3-2=79